# Video LLMs for Temporal Reasoning in Long Videos

### v4

@@ 5.3 Ablation Results
Below we study the effects of our overlapping sliding video Q-Former and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [59].
Impacts of Overlapping Sliding Video Q-Former. We conduct experiments in Tab. 4 to demonstrate the importance of fusing encoded frames across overlapping temporal windows in our clip encoder. Our model with fusion across overlapping windows achieves the best results, whereas using non-overlapping windows or removing the fusion module leads to performance drops. The results validate the design of our overlapping sliding video Q-Former module.
@@
Impacts of BiLSTM. To study the effects of BiLSTM, we replace it with various alternatives and report the results in Tab. 5. From Tab. 5, BiLSTM outperforms average pooling, which is not trained and employed in prior works [29, 50], and traditional LSTM, which only performs the forward pass and hence relies only on past cues. In contrast, BiLSTM is learnable and conducts both forward and backward passes to utilize both past and future cues. Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers [57, 36] alone, leading to worse results. This highlights the importance of BiLSTM even when video Q-Formers [57, 36] are being used. Finally, BiLSTM outperforms transformer [44], which we will discuss below.
@@
Table 5: Effects of the BiLSTM module. Best results are in bold, while second best ones are underlined.
(Bi)LSTM vs. Transformer. The permutation-invariant nature of the self-attention mechanism in transformer [44] leads to loss in temporal information of the input sequence. Though using positional encoding alleviates this issue, it is inevitable that some temporal information loss occurs [56]. In contrast, (Bi)LSTM processes data one by one and preserves the order of the sequence, making them better suited for temporal modeling. Similar to [56], we also observe in Tab. 5 that linear layer outperforms transformer [44].
@@ 5.4 Qualitative Results
Figure 4: Example dense video captioning results in the zero-shot setting on a YouCook2 [59] video. Inaccuracies are shown in red.
Figure 5: Example temporal action segmentation results in the supervised setting on an IndustryASM video. Colored segments represent predicted actions with a particular color denoting the same action across all models. Black represents background frames.
Dense Video Captioning Results. Fig. 4 shows example results of dense video captioning in the zero-shot setting on a YouCook2 [59] video. TimeChat [36] fails to predict any correct timestamps and tends to hallucinate. LongVLM [50] predicts some timestamps correctly and generates more accurate captions, although it still hallucinates objects not present in the video and produces inaccurate/repeated captions. In contrast, TemporalVLM produces much more accurate timestamps and captions. Also, it is the only model that provides captions till the end of the video, showing the effectiveness of the model for long videos.
Temporal Action Segmentation Results. Fig. 5 illustrates example results of temporal action segmentation in the supervised setting on an IndustryASM video. TimeChat [36] falls short in segmenting the video correctly and hallucinates actions not present in the video. LongVLM [50] fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video. In contrast, TemporalVLM predicts action segments significantly closer to the ground truth.
@@ 6 Conclusion
We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos. Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts fused time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features. The extracted features not only are time-sensitive but also contain both local and global cues. In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for time and motion studies and temporal action segmentation evaluation. Lastly, we conduct extensive experiments on TimeIT [36] and IndustryASM to demonstrate our superior results over previous methods across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to incorporate LSTMs into video LLMs.
Appendix A Supplementary Material
In this supplementary material, we first provide additional details of our IndustryASM dataset and our TemporalVLM implementation in Secs. A.1 and A.4 respectively. Next, we provide a comparison with specialized models such as Vid2Seq [53] and QD-DETR [30] in Sec. A.5, followed by demonstrating the effects of number of training epochs in Sec. A.6. Sec. A.7 compares the sizes of our model, TimeChat [36], and LongVLM [50], while Sec. A.8 assesses the impact of number of clips. Sec. A.9 evaluates our model for general video understanding. We then present several qualitative results in Sec. A.10, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results. Finally, we discuss the limitations and societal impacts of our work in Sec. A.11.
A.1 IndustryASM Details
A.2 Dataset Statistics
Our IndustryASM dataset comprises of 4803 videos in total and the average video duration is 119 seconds. Therefore, the total dataset duration is 158 hours. These videos are distributed among 52 industry assembly processes or datasets, ranging from automotive manufacturing, electronic device manufacturing, medical device manufacturing to heating, ventilation, and air conditioning (HVAC) manufacturing. Each assembly process or dataset involves 12 steps or actions on average. We categorize each step into one of the following classes:
@@ A.2 Dataset Statistics
  • Assembling: Here, the worker(s) actively build or complete subassemblies by adding components or combining parts to form a complete unit.
@@
Examples of these 4 categories are presented in Fig. 3 of the main paper. We further classify each assembly process or dataset into one category which appears the most among all of its steps. For example, the Air_Cleaner dataset includes 10 steps (i.e., 4 moving steps, 3 assembling steps, and 3 positioning steps), and hence it is classified as “moving” which appears the most among its 10 steps. We provide the statistics for each category as well as the entire IndustryASM dataset in Tab. 6.
(a) Complete Dataset
Number of videos 4803
Average video duration 119 seconds
Total dataset duration 158 hours
Average number of steps 12
Number of datasets 52
(b) Moving
Number of videos 383
Average video duration 176 seconds
Total dataset duration 19 hours
Average number of steps 11
Number of datasets 7
(c) Assembling
Number of videos 1671
Average video duration 144 seconds
Total dataset duration 67 hours
Average number of steps 14
Number of datasets 20
(d) Positioning
Number of videos 1696
Average video duration 109 seconds
Total dataset duration 51 hours
Average number of steps 13
Number of datasets 18
(e) Packaging
Number of videos 1053
Average video duration 74 seconds
Total dataset duration 22 hours
Average number of steps 10
Number of datasets 7
Table 6: Dataset statistics for IndustryASM.
A.3 Annotation Process
Our IndustryASM videos are annotated with action names and action timestamps by industrial engineers. To ensure quality, two labelers are assigned for each video, i.e., one labeler provides the labels for the video, while another labeler checks the labels. If there are conflicts, both labelers discuss to fix them. In general, roughly 8% of the videos have conflicts and need fixing, resulting in an agreement rate of around 92% between labelers. We manually write the user instructions for the temporal action segmentation task and generate the ground truth responses by using the action names and action timestamps. Fig. 6 illustrates examples of our instruction and ground truth response for the temporal action segmentation on an IndustryASM video.
@@ A.3 Annotation Process
A.4 Implementation Details
We present a detailed list of hyperparameter values used in tuning our TemporalVLM model in Tab. 7. We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames. The frame encoder processes frames of sizes $224\times 224$ across $14\times 14$ patches. Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively. Inside the clip encoder, we divide the timestamp encoded input to the video Q-former into windows of size $q$ and overlap $o$. We then fuse the concatenated window-wise outputs of the Q-former via multiheaded self attention with 8 attention heads that have the same hidden size $N_{v}$ as the Q-Former. The output of the clip encoder contains a fused time-aware representation of each clip. We then concatenate the time-aware clip features according to their temporal order of appearance in the video. This sequence is of dimensions $(C*N^{c}_{f},~{}N_{V})$, with $N_{V}$ denoting the dimension of the video tokens. This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C*N^{c}_{f},~{}2*N_{V})$. The BiLSTM module comprises of two hidden layers, with one hidden layer for the forward LSTM and another hidden layer for the backward LSTM. Projection layers are used to project the BiLSTM output into the LLM input space. The first layer projects the BiLSTM output to the dimensions of $(C*N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
@@ A.4 Implementation Details
Frame encoder patch size 14$\times$14
Frame resolution 224$\times$224
@@
No of attention heads in fusion module 8
@@
A.5 Comparison With Specialized Models
Tab. 8 shows generalist models underperform specialized models such as Vid2Seq  [53] and QD-DETr [30] whose advantages come from task-specific training data and model designs. In particular, the use of 4 loss functions for 200 epochs during training and the utilization of saliency tokens for saliency prediction in QD-DETR  [30] leads to superior performance in highlight detection tasks. In contrast our TemporalVLM is trained on a simpler language modeling loss on lesser epochs. Generalist models such as ours exhibit better generalization across zero-shot, multi-task, and multi-domain settings.
@@ A.5 Comparison With Specialized Models
SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
Vid2Seq 7.9 47.1 27.3
QD-DETR 38.9 62.4
QD-DETR w/ Audio 39.0 62.9
MMN 50.5 29.7
VDI 52.3 31.4
TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 8: Comparisons with specialized models.
A.6 Effects of Number of Training Epochs
Tab. 9 shows the performance of our models across different training epochs.Since our model uses a learnable fusion layer in the clip encoder and a BiLSTM module to aggregate clip level features into a global representation, we observe the best results by training for 7 epochs. However, Tab. 9 reports our results with 3 epochs, which underperform those with 7 epochs but still outperform those of TimeChat and LongVLM.
Model DVC VHD TVG
SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
LongVLM (3 epochs) 2.3 8.1 16.9 16.0 22.5 27.2 11.9
TimeChat (3 epochs) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
TemporalVLM (3 epochs) 3.2 12.9 19.7 23.9 42.3 50.0 25.9
TemporalVLM (7 epochs) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 9: Effects of number of training epochs.
A.7 Comparison on Model Sizes
Tab. 10 compares the sizes of our TemporalVLM model, TimeChat [36], and LongVLM [50] in terms of the number of learnable parameters. Our TemporalVLM model includes a trainable BiLSTM module and projection layers, leading to a 5% and 4% increase in the number of trainable parameters over TimeChat [36] and LongVLM [50] respectively. Nevertheless, our TemporalVLM model achieves the best performance across various temporal reasoning and fine-grained understanding tasks, despite using less training data than TimeChat [36]. In particular, TimeChat [36] is trained on the complete TimeIT [36] and Valley [28] datasets, whereas our TemporalVLM model is trained on a subset [^2] We could not download the YT-Temporal [55] dataset on time due to its large size and the restricted number of downloads by YouTube. of the TimeIT [36] and Valley [28] datasets.
A.8 Impacts of Number of Short-Term Clips
Tab. 11 presents the effects of the number of short-term clips on our TemporalVLM model. Due to memory limitations, we have experimented with three values for the number of short-term clips, namely 2, 4, and 6. It is evident from Tab. 11 that the performance is improved with increasing the number of short-term clips, since our model is able to access more data from the input video.
Model Number of Trainable Parameters
TimeChat [36] 241 millions
LongVLM [50] 244 millions
TemporalVLM (Ours) 255 millions
Table 10: Comparison on model sizes.
# Clips SODA_c CIDEr F1

### v5

@@ 5.3 Ablation Results
We fine-tune several variants of TemporalVLM on YouCook2 Zhou et al. (2018) to study our overlapping sliding video Q-Former and BiLSTM.
Impacts of Overlapping Sliding Video Q-Former. Tab. 4 demonstrates the importance of fusing encoded frames across overlapping temporal windows in our clip encoder. Our model with fusion across overlapping windows achieves the best results, whereas using non-overlapping windows and removing the fusion module as in TimeChat’s Ren et al. (2024) design lead to performance drops. The results validate our design.
@@
Impacts of BiLSTM. To study the effects of BiLSTM, we replace it with various alternatives and report the results in Tab. 5. From Tab. 5, BiLSTM outperforms average pooling, which is not trained and employed in prior works Maaz et al. (2023); Weng et al. (2024), and traditional LSTM, which only performs the forward pass and hence relies only on past cues. In contrast, BiLSTM is learnable and conducts both forward and backward passes to utilize both past and future cues. Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers Zhang et al. (2023); Ren et al. (2024) alone, leading to worse results. This highlights the importance of BiLSTM even when video Q-Formers are being used. Finally, BiLSTM outperforms transformer Vaswani et al. (2017), which we will discuss below.
@@
Table 5: Effects of BiLSTM module. Best results are in bold, while second best ones are underlined.
(Bi)LSTM vs. Transformer. The permutation-invariant nature of the self-attention mechanism in transformer Vaswani et al. (2017) leads to loss in temporal information of the input sequence. Though using positional encoding alleviates this issue, it is inevitable that some temporal information loss occurs Zeng et al. (2023). In contrast, (Bi)LSTM processes data one by one and preserves the sequence order, making them better suited for temporal modeling. Similar to Zeng et al. (2023), we also observe in Tab. 5 that linear layer outperforms transformer Vaswani et al. (2017).
@@ 5.4 Qualitative Results
Figure 4: Dense video captioning in zero-shot setting on YouCook2. Red denotes inaccuracies.
Figure 5: Temporal action segmentation in supervised setting on IndustryASM. Black denotes background.
Dense Video Captioning Results. Fig. 4 shows example results in the zero-shot setting on a YouCook2 Zhou et al. (2018) video. TimeChat Ren et al. (2024) fails to predict any correct timestamps and tends to hallucinate. LongVLM Weng et al. (2024) yields better results but still hallucinates objects not present in the video and produces inaccurate captions. In contrast, TemporalVLM produces much more accurate timestamps and captions. Also, it is the only model that provides captions till the end of the video, showing the effectiveness for long video understanding.
Temporal Action Segmentation Results. Fig. 5 presents example results in the supervised setting on an IndustryASM video. TimeChat Ren et al. (2024) fails to segment the video and hallucinates actions not present in the video. LongVLM Weng et al. (2024) fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video. In contrast, TemporalVLM predicts the action segments notably closer to the ground truth.
Supplementary Material. Due to space constraints, we include additional details, evaluations, and discussions in our supplementary material.
@@ 6 Conclusion
We propose TemporalVLM for temporal reasoning and fine-grained understanding in long videos. Our approach includes a time-aware clip encoder for extracting fused time-aware local features and a BiLSTM module for global feature aggregation. Extracted features not only are time-sensitive but also contain both local and global cues. Moreover, we present IndustryASM, a long video dataset of industry assembly processes. Lastly, extensive experiments demonstrate our superior results over prior works across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to blend LSTMs into video LLMs. Our future work will explore advanced recurring models Gu et al. (2021); Gu and Dao (2024).
Supplementary Material
In this supplementary material, we first provide the additional details of our IndustryASM dataset (including dataset statistics, annotation process, maintenance and release plans) and our TemporalVLM implementation in Secs. A and B respectively. Next, we provide the comparisons with specialized models (e.g., Vid2Seq Yang et al. (2023) and QD-DETR Moon et al. (2023)) in Sec. C, followed by demonstrating the effects of number of training epochs in Sec. D. We then evaluate our model for general video understanding in Sec. E. Furthermore, Sec. F compares the sizes of our model against TimeChat Ren et al. (2024) and LongVLM Weng et al. (2024), while we report the run time comparisons in Sec. G to evaluate the efficiency of our model. Moreover, Sec. H assesses the impacts of number of clips, while we study the effects of employing a more advanced LLM in our model in Sec. I. We then present several qualitative results in Sec. J, including dense video captioning, video highlight detection, temporal video grounding, and more importantly, generalization results. Finally, we discuss the limitations and societal impacts of our work in Secs. K and L respectively.
Appendix A IndustryASM Details
A.1 Dataset Statistics
Our IndustryASM dataset comprises of 4803 videos in total and the average video duration is 119 seconds. Therefore, the total dataset duration is 158 hours. These videos are distributed among 52 industry assembly processes or datasets, ranging from automotive manufacturing, electronic device manufacturing, medical device manufacturing to heating, ventilation, and air conditioning (HVAC) manufacturing. Participants in our dataset are salaried factory workers based in the US. Informed consent is obtained from all participants in accordance with standard operating procedures prior to data collection. As an additional privacy safeguard, all faces and other personally identifiable information (PII) are blurred or otherwise anonymized in the recorded data. Each assembly process or dataset involves 12 steps or actions on average. We first categorize each step into one of the following classes:
@@ A.1 Dataset Statistics
  • Assembling: The worker(s) build or complete subassemblies by adding components or combining parts to form a complete unit.
@@
Next, we further classify each assembly process or dataset into one of the above categories which appears the most among all of its steps. For example, the Air_Cleaner dataset includes 10 steps (namely, 4 moving steps, 3 assembling steps, and 3 positioning steps), and hence it is classified as “Moving” which appears the most among its 10 steps. We provide the statistics for each category of the IndustryASM dataset in Fig. 6.
Finally, examples of these 4 categories are presented in Fig. 3 of the main paper. In addition, Fig. 7 presents more examples to better illustrate the diverse camera viewpoints, actors, backgrounds, and activities in our IndustryASM dataset that are common in manufacturing settings.
Figure 6: Dataset statistics for our IndustryASM dataset.
@@
A.2 Annotation Process
Our IndustryASM videos are annotated with action names and action timestamps by industrial engineers. To ensure quality, two labelers are assigned for each video, i.e., one labeler provides the labels for the video, while another labeler checks the labels. If there are conflicts, both labelers discuss to fix them. In general, roughly 8% of the videos have conflicts and need fixing, resulting in an agreement rate of around 92% between labelers. We manually write the user instructions for the temporal action segmentation task and generate the ground truth responses by using the action names and action timestamps. Fig. 8 illustrates examples of our instruction and ground truth response for temporal action segmentation on an IndustryASM video.
A.3 Maintenance and Release Plans
We will host our IndustryASM dataset on Hugging Face and assign a permanent DOI to it to ensure long-term accessibility. We will release version updates to incorporate any corrections and add additional data/annotations if and when available. These updates will follow a versioning scheme (e.g., v1.0) and there will be documentations on any changes in the release notes. We will track and respond to user reported issues and errors on our GitHub repository. We will release the entire dataset upon acceptance of this paper and we plan to actively maintain it for at least 5 years following the release date, after which the dataset will remain archived with its DOI.
Appendix B Implementation Details
We perform grid search to find the best hyperparameter settings for TemporalVLM on supervised dense video captioning on YouCook2 Zhou et al. (2018) and use them for all the remaining experiments. We present our hyperparameter settings in Tab. 6. We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames. The frame encoder processes frames of sizes $224\times 224$ across $14\times 14$ patches. Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively. Inside the clip encoder, we divide the timestamp encoded input to the video Q-former into windows of size $q=32$ frames and overlap $o=16$ frames. We then fuse the concatenated window-wise outputs of the Q-former via multi-headed self-attention with 8 attention heads that have the same hidden size $N_{v}$ as the Q-Former. The output of the clip encoder contains a fused time-aware representation of each clip.
We then concatenate the time-aware clip features according to their temporal order of appearance in the video. This sequence is of dimensions $(C\times N^{c}_{f},~N_{V})$, with $N_{V}$ denoting the dimension of the video tokens. This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C\times N^{c}_{f},~2\times N_{V})$. The BiLSTM module comprises of two hidden layers, with one hidden layer for the forward LSTM and another hidden layer for the backward LSTM. Projection layers are used to project the BiLSTM output into the LLM input space. The first layer projects the BiLSTM output to the dimensions of $(C\times N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
@@ Appendix B Implementation Details
Frame encoder patch size $\times14$
Frame resolution $\times224$
@@
Number of attention heads in fusion module 8
@@
Appendix C Comparisons with Specialized Models
Tab. 7 shows that specialized models such as Vid2Seq Yang et al. (2023) and QD-DETR Moon et al. (2023), whose advantages come from task-specific training data and model designs, outperform generalist models such as TimeChat Ren et al. (2024) and our TemporalVLM model. In particular, the use of 4 loss functions for 200 epochs during training and the utilization of saliency tokens for saliency prediction in QD-DETR lead to superior performance in video highlight detection tasks. In contrast, our TemporalVLM model is trained on a simpler language modeling loss for a smaller number of epochs. Nevertheless, generalist models exhibit better generalization across zero-shot, multi-task, and multi-domain settings.
Model Dense Captioning Highlight Detection Temporal Grounding
YouCook2 QVHighlights Charades-STA
SODA_c CIDEr F1 mAP HIT@1 R@1 (IoU=0.5) R@1 (IoU=0.7)
Vid2Seq Yang et al. (2023) 7.9 47.1 27.3
QD-DETR Moon et al. (2023) 38.9 62.4
QD-DETR w/ Audio Moon et al. (2023) 39.0 62.9
MMN Wang et al. (2022) 50.5 29.7
VDI Luo et al. (2023a) 52.3 31.4
Timechat Ren et al. (2024) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 7: Comparisons with specialized models. Best results are in bold, while second best ones are underlined.
Appendix D Effects of Number of Training Epochs
Tab. 8 shows the performance of our model across different training epochs. Since our model uses a learnable fusion module in the clip encoder and a learnable BiLSTM module to aggregate clip-level features into a global representation, we observe the best results by training for 7 epochs. Tab. 8 reports our results with 3 epochs, which are worse than ours with 7 epochs but still outperform those of TimeChat Ren et al. (2024) and LongVLM Weng et al. (2024).
Model Epochs Dense Captioning Highlight Detection Temporal Grounding
YouCook2 QVHighlights Charades-STA
SODA_c CIDEr F1 mAP HIT@1 R@1 (IoU=0.5) R@1 (IoU=0.7)
TimeChat Ren et al. (2024) 3 3.1 10.3 19.5 21.7 37.9 46.7 23.7
LongVLM Weng et al. (2024) 3 2.3 8.1 16.9 16.0 22.5 27.2 11.9
TemporalVLM 3 3.2 12.9 19.7 23.9 42.3 50.0 25.9
TemporalVLM 7 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 8: Effects of number of training epochs. Best results are in bold, while second best ones are underlined.
Appendix E General Video Understanding Results
Following LongVLM Weng et al. (2024), we evaluate the performance of our TemporalVLM model on the general video understanding benchmark provided by Video-ChatGPT Maaz et al. (2023) in Tab. 9. The evaluation metrics include Correctness Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (C). It is evident from Tab. 9 that our TemporalVLM model achieves the best performance across all metrics, outperforming all competing methods, including the original LongVLM model.
Model Data CI DO CU TU C
Video-LLaMA Zhang et al. (2023) 10M 1.96 2.18 2.16 1.82 1.79
Video-ChatGPT Ren et al. (2024) 100K 2.50 2.57 2.69 2.16 2.20
Valley Luo et al. (2023b) 234K 2.43 2.13 2.86 2.04 2.45
BT-Adapter Liu et al. (2024) 10M 2.68 2.69 3.27 2.34 2.46
LongVLM Weng et al. (2024) 100K 2.76 2.86 3.34 2.39 3.11
TemporalVLM 100K 2.88 2.91 3.45 2.50 3.16
Appendix F Model Size Comparisons
Tab. 10 compares the sizes of our TemporalVLM model, TimeChat Ren et al. (2024), and LongVLM Weng et al. (2024) in terms of the number of learnable parameters (measured in millions). Our TemporalVLM model includes a learnable fusion module and a learnable BiLSTM module, leading to a 5% and 4% increase in the number of trainable parameters over TimeChat and LongVLM respectively. Nevertheless, our TemporalVLM model achieves the best performance across various temporal reasoning and fine-grained understanding tasks, despite using less training data than TimeChat. In particular, TimeChat is trained on the complete TimeIT Ren et al. (2024) and Valley Luo et al. (2023b) datasets, whereas our TemporalVLM model is trained on a subset [^2] We could not download YT-Temporal Zellers et al. (2022) due to its large size and the restricted number of downloads. of the TimeIT and Valley datasets.
Model Parameters
TimeChat Ren et al. (2024) 241M
LongVLM Weng et al. (2024) 244M
TemporalVLM 255M
Table 10: Model size comparisons.
Appendix G Run Time Comparisons
We compare the inference times (measured in seconds) of different models on three tasks, i.e., Dense Video Captioning (DVC) on YouCook2 Zhou et al. (2018), Video Highlight Detection (VHD) on QVHighlights Lei et al. (2021), and Temporal Video Grounding (TVG) on Charades Gao et al. (2017) in Tab. 11. From the results, TimeChat Ren et al. (2024) is about 35% more efficient than LongVLM Weng et al. (2024) and TemporalVLM, since both LongVLM and TemporalVLM perform clip sampling and clip encoding.
@@ Appendix G Run Time Comparisons
TimeChat Ren et al. (2024) 10.1s 5.6s 3.7s
LongVLM Weng et al. (2024) 13.2s 7.3s 4.9s
TemporalVLM 13.5s 7.6s 5.1s
Table 11: Run time comparisons.
Appendix H Effects of Number of Short-Term Clips
Tab. 12 presents the effects of number of short-term clips on our TemporalVLM model. Due to memory limitations, we have experimented with three values for the number of short-term clips, namely 2, 4, and 6. It is evident from Tab. 12 that the performance is improved with increasing the number of short-term clips, since our model is able to access more data from the input video.
Clips SODA_c CIDEr F1